The speed of a train is reduced uniformly from 15 m/s to 7 m/s while traveling a distance of 90 m. How much farther will the train travel before coming to rest, provided the acceleration remains constant? The acceleration is -9.81 m/s2


1
Expert's answer
2021-07-19T09:51:40-0400

"v^2+u^2=2as \\

u= 15 \;m/s \\

v=7 \;m/s \\

a=90 \;m \\

15^2 + 7^2 = 2a \times 90 \\

\frac{176}{180} = a \\

a=0.98 \;m/s^2"

Now the final speed "v_f=0"

"u=7\;m/s \\

v_f^2+u^2=2as \\

0 -7^2 = 2 \times 0.98 s \\

s = \frac{49}{1.96} \\

s= 25 \;m"